With block C placed on top of B. the system undergoes simple harmonic motion with an amplitude of 0.12 m. Block B has a speed of 0.22 m/s at a displacement of 0.075 m from its equilibrium position. Determine the period of the motion. Express your answer with the appropriate units. What minimum value for the coefficient of static friction mu_s between B and C is needed if C is never to slip? Ignore any friction between B and the horizontal surface.


1
Expert's answer
2021-08-16T08:34:11-0400

We know the amplitude, so, we know the total energy of the system:


"E_T=\frac12 kA^2."

This is equivalent to the sum of elastic potential and kinetic energy of the system at the moment when the speed is 0.22 m/s and displacement of 0.075 m.


"E_T=\frac12 kx^2+\frac12Mv^2."

The period of motion is


"T=2\pi\sqrt{\frac{M}{k}}."

So, what we have is


"\frac12 kA^2=\frac12 kx^2+\frac12Mv^2,\\\space\\
k=\frac{Mv^2}{A^2+x^2}."

Substitute this into the equation for period:


"T=\frac{2\pi}{v}\sqrt{A^2+x^2}=4\text{ s}."

Now, let's talk about the coefficient of friction. Write the equation for displacement:


"x(t)=A\sin\bigg(\frac{2\pi}{T}t\bigg),"

the acceleration is the second derivative of displacement by time:


"-a(t)=A\bigg(\frac{2\pi}{T}\bigg)^2\sin\bigg(\frac{2\pi}{T}t\bigg)."

So, the magnitude of acceleration is


"|a|=A\bigg(\frac{2\pi}{T}\bigg)^2=0.3\text{ m/s}^2."

The coefficient of friction is


"\mu_s=\frac ag=0.03."